# When are off-diagonal hypergraph Ramsey numbers polynomial?

### v2

@@ 2 Tightly connected hypergraphs
Let $\ell=C\log N$, where $C>0$ is a sufficiently large absolute constant, and $n=\ell^{3/2}$. Let $r=0.01\ell$ and let $V$ be an $r$-trifference code in $\{1,2,3\}^{\ell}$ of size $N$. What this means is that $V$ is a subset of $\{1,2,3\}^{\ell}$ of size $N$ with the property that, for any triple $uvw$ of elements of $V$, there are at least $r$ coordinates where $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$. Such a set of size $N$ can be found by the first moment method, as it is exponentially unlikely that a random triple $uvw$ is not $r$-trifferent.
We will define our Ramsey coloring on the complete $3$-graph with vertex set $V$. Let $c(uv)$ denote the set of coordinates in which $u,v$ differ. By the definition of $V$, $r\leq|c(uv)|\leq\ell$ for all $u,v$. Define $\phi(uv)$ to be an element of $c(uv)$ picked uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if $\phi(uv)=\phi(vw)=\phi(uw)$. Observe that, by definition, any red triple $uvw$ automatically satisfies $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$ for $i=\phi(uv)$.
@@
No blue $K_{n}^{(3)}$ with positive probability. Fix a set $U$ of $n$ vertices in $V$; we would like to bound the probability that this $n$-set forms a blue clique. Let $T_{uvw}$ be the event that the triple $uvw$ is red, that is, that the triangle $uvw$ is monochromatic under $\phi$. Each event has probability
@@
where $\tau$ ranges over all triangles $uvw$ in $U$ and $\tau\sim\tau^{\prime}$ if these two triangles share at least one edge, with
@@
Before moving on to the case with two tight components, we sketch an alternative, but quantitatively weaker, construction for the single component case. Let $\ell=C\log N$ and $n=C^{7/4}\log^{2}N$ for $C$ a sufficiently large constant. For each edge $uv$ of the complete graph on $V=[N]$, label it with some $\phi(uv)\in\{1,2,...,\ell\}$ chosen independently and uniformly at random. Moreover, assign a string $f(v)\in\{1,2,3\}^{\ell}$ to each vertex $v$, again chosen independently and uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if
@@
and blue otherwise. If there is any red tightly connected $3$-graph in $\chi$, then all edges in its $2$-shadow have the same $\phi$-value $i$, so the $f_{i}(v)$ yield a tripartition. To show that there is no blue $K_{n}$ with positive probability, we use an application of Janson’s inequality to show that, with positive probability, for every subset $U$ of $V$ with $n$ vertices and every $i\in\{1,2,...,\ell\}$ there are at least $n/6$ vertex-disjoint triangles in $U$ which are monochromatic of color $i$ under $\phi$. If we now bring the randomness of $f$ into play, then each such monochromatic triangle forms a blue edge with probability $7/9$. As the randomness for each coordinate $f_{i}$ of $f$ is independent, the induced hypergraph on any given $n$-vertex set $U$ is completely blue with probability at most $(7/9)^{\frac{n}{6}\cdot\ell}=\exp(-\Theta(n\ell))$. Since there are (${N\choose n}=\exp(O(n\log N))$ subsets to consider, the union bound and the fact that $\ell$ dominates $\log N$ completes the analysis, yielding the bound $r(H,K_{n}^{(3)})\geq 2^{\Omega(n^{1/2})}$ for every non-tripartite tightly connected $3$-graph $H$.
@@ 3 Two tight components
Replace $N$ by the smallest power of $2$ exceeding $N$ and let $\ell=\log_{2}N$. Let $\varphi$ be as in Lemma 3.2 with $A,k=\exp(\Theta(\sqrt{\ell}))$. Let $g,f_{1},f_{2},f_{3}$ be four independent random functions on (${[N]\choose 2}$, where $g:{[N]\choose 2}\rightarrow\{1,2,3\}$ uniformly at random and $f_{1},f_{2},f_{3}:{[N]\choose 2}\rightarrow[\ell A]$ uniformly at random. The Ramsey coloring of the complete $3$-graph on $[N]$ will be $\chi$ where $\chi(uvw)$ for a triple $uvw$ with $u<v<w$ is red if and only if
@@
    $uvw$ is a rainbow triangle with respect to $\varphi$,
@@
Otherwise, the triple $uvw$ is colored blue.
@@
Suppose now that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are two red tight components whose union is not tripartite. Without loss of generality, this means that there is some $u\in V^{(1)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$ and $v\in V^{(2)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$. If some $w\in V^{(3)}_{\mathcal{C}}$ is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}^{\prime}}$ and $V^{(i)}_{\mathcal{C}^{\prime}}$ has the same color under $\varphi$ for each of $i=2,3$, $c^{(2)}_{\mathcal{C}}=\varphi(uw)=\varphi(vw)=c^{(1)}_{\mathcal{C}}$, which is a contradiction. Therefore, $V^{(3)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. If there is some $w\in V^{(1)}_{\mathcal{C}}$ that is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, say $w\in V^{(2)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}}$ and $V^{(2)}_{\mathcal{C}}$ has the same color under $\varphi$, $\varphi(uv)=c^{(3)}_{\mathcal{C}}=\varphi(vw)=c^{(3)}_{\mathcal{C}^{\prime}}$. But then picking any $x\in V^{(2)}_{\mathcal{C}^{\prime}}$ we have a monochromatic triangle $uvx$, which contradicts the fact that each color class in $\varphi$ is bipartite. Therefore, $V^{(1)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must also be empty. Similarly, $V^{(2)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. Hence, $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$ and $V_{\mathcal{C}}$ are disjoint, so the union of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ is iterated tripartite, as required. ∎
@@
$(t,t^{\prime})=\begin{cases}(u,v)&\textup{ if }\frac{m_{u}}{n_{u}-2}\geq\frac{m_{v}}{n_{v}-2},\\ (v,u)&\textup{ otherwise},\end{cases}$
then $M(W^{\prime})$ is a line through the origin with slope $m_{t}/(n_{t}-2)$ on $[0,n_{t}-2]$ and the slope becomes $m_{t^{\prime}}/(n_{t^{\prime}}-2)$ on $[n_{t}-2,n_{u}+n_{v}-4]$.
@@
so $M^{\prime}(W^{\prime})$ is a line through the origin with slope $m_{u}^{\prime}/(n_{u}^{\prime}-2)$ on $[0,n_{u}^{\prime}-2]$ and the slope becomes $m_{v}^{\prime}/(n_{v}^{\prime}-2)$ on $[n^{\prime}_{u}-2,n_{u}^{\prime}+n_{v}^{\prime}-4]$, after which it remains constant.
@@
We may therefore assume that $T$ has no imbalances. In particular, for any vertex $u$ with children $v$ and $w$, we must have that $n_{v},n_{w}$ are both in $[n_{u}/3,2n_{u}/3]$. Otherwise, without loss of generality, we may assume that $n_{w}<n_{u}/3$ and $n_{v}>2n_{u}/3$. But then there would exist a child $x$ of $v$ such that $n_{x}>n_{u}/3>n_{w}$, contradicting that there are no imbalances. Therefore, any vertex $u$ at depth $d<0.1\log k$ is an internal vertex with weight $\Omega(3^{-d}k)$ and $m_{u}=\Omega(3^{-3d}k^{3})$. Since $\Omega(3^{-3d}k^{3})=\omega(k^{2.5})$, if the score of a function $f:V_{\textup{in}}\to[0,1]$ is less than $k^{2.5}$, then any vertex whose depth is less than $0.1\log k$ must be mapped to a number which is at least $1/2$ by $f$. Since the vertices at depth $d$ have total weight $k-O(2^{d})$ for every $d<0.1\log k$, it follows that the weight of $f$ is $\Omega(k\log k)$, as desired. ∎
@@
If $S^{\prime}$ is good, then we know that, for every $u\in S_{0}^{\prime}$, there are at least $\left\lvert S_{0}^{\prime}\right\rvert/2$ vertices $v\in S_{0}^{\prime}$ with $c_{t}(u)\neq c_{t}(v)$. For any such $u,v\in S_{0}^{\prime}$ and any $w\in S_{1}^{\prime}$, it is clear that $uvw$ is rainbow. Therefore, we have found at least
$\frac{1}{2}\frac{\left\lvert S_{0}^{\prime}\right\rvert^{2}}{2}\left\lvert S_{1}^{\prime}\right\rvert\geq\frac{1}{4}\left(\binom{\left\lvert S_{0}^{\prime}\right\rvert}{2}\left\lvert S_{1}^{\prime}\right\rvert+\binom{\left\lvert S_{1}^{\prime}\right\rvert}{2}\left\lvert S_{0}^{\prime}\right\rvert\right)$
@@
If we now apply Lemma 3.4 with $X$ the set of bad vertices in $T$, we see that if the total weight on the vertices in $X$ is less than $k\log k/C$, then the number of triples of leaves whose least common ancestor is outside $X$, and therefore good, is at least $k^{2.5}$. By the observation above, this would mean that we have at least $k^{2.5}/4$ rainbow triangles in $S$. We may therefore assume that the total weight on the set of bad vertices is at least $k\log k/C$. Since the weight of the vertex corresponding to $S^{\prime}$ is $\left\lvert S^{\prime}\right\rvert-2$ and the probability that $S^{\prime}$ is bad can be bounded by $\exp(-\Omega((\left\lvert S^{\prime}\right\rvert-2))\log A)$, the probability that all of the vertices in $X$ are bad is at most $\exp(-\Omega(k\log k\log A))$. Since there are at most $2^{k}$ possible choices for $X$ (as there are exactly $k-1$ internal vertices), a union bound implies that the probability there are fewer than $k^{2.5}/4$ rainbow triangles in $S$ is at most $\exp(-\Omega(k\log k\log A))$.
@@ 4 Concluding remarks
To prove the claim, we pick an arbitrary vertex $v\in V(K_{n}^{(3)})$ and consider $G_{v}$, the red/blue-colored graph on $V\setminus\{v\}$ where the color of each edge $wx$ is the color of the triple $vwx$. Since, by assumption, all red tight components in our coloring of $K_{n}^{(3)}$ are tripartite, all red connected components of $G_{v}$ are bipartite. This is exactly equivalent to saying that the set of red edges in $G_{v}$ is bipartite. Therefore, there is a subset $V^{\prime}$ of $V\backslash\{v\}$ such that $G_{v}[V^{\prime}]$ contains only blue edges and $\left\lvert V^{\prime}\right\rvert\geq\lceil\frac{N-1}{2}\rceil$. Now consider the coloring of the $3$-graph on $V^{\prime}$, which again has no red copy of any hypergraph from $\mathcal{H}$. By the definition of the function $n$, there exists a subset $V^{\prime\prime}$ of $V^{\prime}$ of size $n\left(\lceil\frac{N-1}{2}\rceil\right)$ such that $V^{\prime\prime}$ is completely blue. It is then easy to see that $V^{\prime\prime}\cup\{v\}$ is completely blue as well, as the presence of any red edge would contradict the choice of $V^{\prime}$. The required bound on $n(N)$ follows.
@@
Our methods prove this conjecture for $s=2$ and 4.2 follows immediately from 4.3 as $r(H,K_{n}^{(k)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any $H\in\mathcal{H}_{s}^{(k)}$. We note that all the proofs in this paper followed this strategy. Moreover, we observe that 4.3, if true in general, would also determine the tower height of diagonal hypergraph Ramsey numbers, perhaps the central open problem in hypergraph Ramsey theory. Indeed, $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any positive integers $s,k,n$ with $k>s$, so we would have that the tower height of $r(K_{n}^{(s)},K_{n}^{(s)})$ is at least $s-1$, agreeing with the longstanding upper bound.
To see the inequality $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$, suppose for the sake of contradiction that there is a red/blue-coloring of the edges of $K_{N}^{(k)}$ with $N=r(K_{n}^{(s)},K_{n}^{(s)})$ that avoids any red $H$ in $\mathcal{H}_{s}^{(k)}$. Label the $k$ parts of each of the red $s$-tight components from $1$ to $k$. Consider the auxiliary coloring of the edges of $K_{N}^{(s)}$ where an edge $e$ is red if it is contained in a red edge $\tilde{e}$ in the original coloring and $e$ contains precisely the vertices in $\tilde{e}$ from parts $1,2,\ldots,s$ in the red $s$-tight component of $\tilde{e}$; $e$ is blue otherwise. Since $\left\lvert e\right\rvert=s$, any red $\tilde{e}$ containing $e$ belongs to the same red $s$-tight component, showing that the color of $e$ does not depend on the choice of $\tilde{e}$. It is clear that each red $\tilde{e}$ in $K_{N}^{(k)}$ contains exactly one red $e$ in the auxiliary coloring of $K_{N}^{(s)}$. By the definition of $r(K_{n}^{(s)},K_{n}^{(s)})$, this auxiliary coloring contains either a red $K_{n}^{(s)}$ or a blue $K_{n}^{(s)}$. We will show that either one would give a blue $K_{n}^{(k)}$ in the original coloring. Indeed, suppose that $S$ is the set of $n$ vertices that forms a monochromatic $K_{n}^{(s)}$ in the auxiliary coloring. Then each $\tilde{e}\in\binom{S}{k}$ contains either 00 or $\binom{k}{s}>1$ red edges $e$, so it has to be blue. This yields the required contradiction and so $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ must hold.

### v3

@@ 2 Tightly connected hypergraphs
Let $\ell=C\log N$, where $C>0$ is a sufficiently large absolute constant, and $n=\ell^{3/2}$. Let $r=0.01\ell$ and let $V$ be an $r$-trifference code in $\{1,2,3\}^{\ell}$ of size $N$. What this means is that $V$ is a subset of $\{1,2,3\}^{\ell}$ of size $N$ with the property that, for any triple uvwuvw of elements of $V$, there are at least $r$ coordinates where $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$. Such a set of size $N$ can be found by the first moment method, as it is exponentially unlikely that a random triple uvwuvw is not $r$-trifferent.
We will define our Ramsey coloring on the complete $3$-graph with vertex set $V$. Let $c(uv)$ denote the set of coordinates in which $u,v$ differ. By the definition of $V$, $r\leq|c(uv)|\leq\ell$ for all $u,v$. Define $\phi(uv)$ to be an element of $c(uv)$ picked uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if $\phi(uv)=\phi(vw)=\phi(uw)$. Observe that, by definition, any red triple uvwuvw automatically satisfies $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$ for $i=\phi(uv)$.
@@
No blue $K_{n}^{(3)}$ with positive probability. Fix a set $U$ of $n$ vertices in $V$; we would like to bound the probability that this $n$-set forms a blue clique. Let $T_{uvw}$ be the event that the triple uvwuvw is red, that is, that the triangle uvwuvw is monochromatic under $\phi$. Each event has probability
@@
where $\tau$ ranges over all triangles uvwuvw in $U$ and $\tau\sim\tau^{\prime}$ if these two triangles share at least one edge, with
@@
Before moving on to the case with two tight components, we sketch an alternative, but quantitatively weaker, construction for the single component case. Let $\ell=C\log N$ and $n=C^{7/4}\log^{2}N$ for $C$ a sufficiently large constant. For each edge $uv$ of the complete graph on $V=[N]$, label it with some $\phi(uv)\in\{1,2,$$\ell\}$ chosen independently and uniformly at random. Moreover, assign a string $f(v)\in\{1,2,3\}^{\ell}$ to each vertex $v$, again chosen independently and uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if
@@
and blue otherwise. If there is any red tightly connected $3$-graph in $\chi$, then all edges in its $2$-shadow have the same $\phi$-value $i$, so the $f_{i}(v)$ yield a tripartition. To show that there is no blue $K_{n}$ with positive probability, we use an application of Janson’s inequality to show that, with positive probability, for every subset $U$ of $V$ with $n$ vertices and every $i\in\{1,2,$$\ell\}$ there are at least $n/6$ vertex-disjoint triangles in $U$ which are monochromatic of color $i$ under $\phi$. If we now bring the randomness of $f$ into play, then each such monochromatic triangle forms a blue edge with probability $7/9$. As the randomness for each coordinate $f_{i}$ of $f$ is independent, the induced hypergraph on any given $n$-vertex set $U$ is completely blue with probability at most $(7/9)^{\frac{n}{6}\cdot\ell}=\exp(-\Theta(n\ell))$. Since there are ${N\choose n}=\exp(O(n\log N))$ subsets to consider, the union bound and the fact that $\ell$ dominates $\log N$ completes the analysis, yielding the bound $r(H,K_{n}^{(3)})\geq 2^{\Omega(n^{1/2})}$ for every non-tripartite tightly connected $3$-graph $H$.
@@ 3 Two tight components
Replace $N$ by the smallest power of $2$ exceeding $N$ and let $\ell=\log_{2}N$. Let $\varphi$ be as in Lemma 3.2 with $A,k=\exp(\Theta(\sqrt{\ell}))$. Let $g,f_{1},f_{2},f_{3}$ be four independent random functions on $N]\choose 2$}, where $g:{[N]\choose 2}\rightarrow\{1,2,3\}$ uniformly at random and $f_{1},f_{2},f_{3}:{[N]\choose 2}\rightarrow[\ell A]$ uniformly at random. The Ramsey coloring of the complete $3$-graph on $[N]$ will be $\chi$ where $\chi(uvw)$ for a triple uvwuvw with $u<v<w$ is red if and only if
@@
    uvwuvw is a rainbow triangle with respect to $\varphi$,
@@
Otherwise, the triple uvwuvw is colored blue.
@@
Suppose now that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are two red tight components whose union is not tripartite. Without loss of generality, this means that there is some $u\in V^{(1)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$ and $v\in V^{(2)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$. If some $w\in V^{(3)}_{\mathcal{C}}$ is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}^{\prime}}$ and $V^{(i)}_{\mathcal{C}^{\prime}}$ has the same color under $\varphi$ for each of $i=2,3$, $c^{(2)}_{\mathcal{C}}=\varphi(uw)=\varphi(vw)=c^{(1)}_{\mathcal{C}}$, which is a contradiction. Therefore, $V^{(3)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. If there is some $w\in V^{(1)}_{\mathcal{C}}$ that is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, say $w\in V^{(2)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}}$ and $V^{(2)}_{\mathcal{C}}$ has the same color under $\varphi$, $\varphi(uv)=c^{(3)}_{\mathcal{C}}=\varphi(vw)=c^{(3)}_{\mathcal{C}^{\prime}}$. But then picking any $x\in V^{(2)}_{\mathcal{C}^{\prime}}$ we have a monochromatic triangle uvxuvx, which contradicts the fact that each color class in $\varphi$ is bipartite. Therefore, $V^{(1)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must also be empty. Similarly, $V^{(2)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. Hence, $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$ and $V_{\mathcal{C}}$ are disjoint, so the union of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ is iterated tripartite, as required. ∎
@@
[…]
then $M(W^{\prime})$ is a line through the origin with slope $m_{t}/(n_{t}-2)$ on $0,n_{t}-2]$ and the slope becomes $m_{t^{\prime}}/(n_{t^{\prime}}-2)$ on $n_{t}-2,n_{u}+n_{v}-4]$.
@@
so $M^{\prime}(W^{\prime})$ is a line through the origin with slope $m_{u}^{\prime}/(n_{u}^{\prime}-2)$ on $0,n_{u}^{\prime}-2]$ and the slope becomes $m_{v}^{\prime}/(n_{v}^{\prime}-2)$ on $n^{\prime}_{u}-2,n_{u}^{\prime}+n_{v}^{\prime}-4]$, after which it remains constant.
@@
We may therefore assume that $T$ has no imbalances. In particular, for any vertex $u$ with children $v$ and $w$, we must have that $n_{v},n_{w}$ are both in $n_{u}/3,2n_{u}/3]$. Otherwise, without loss of generality, we may assume that $n_{w}<n_{u}/3$ and $n_{v}>2n_{u}/3$. But then there would exist a child $x$ of $v$ such that $n_{x}>n_{u}/3>n_{w}$, contradicting that there are no imbalances. Therefore, any vertex $u$ at depth $d<0.1\log k$ is an internal vertex with weight $\Omega(3^{-d}k)$ and $m_{u}=\Omega(3^{-3d}k^{3})$. Since $\Omega(3^{-3d}k^{3})=\omega(k^{2.5})$, if the score of a function $f:V_{\textup{in}}\to[0,1]$ is less than $k^{2.5}$, then any vertex whose depth is less than $0.1\log k$ must be mapped to a number which is at least $1/2$ by $f$. Since the vertices at depth $d$ have total weight $k-O(2^{d})$ for every $d<0.1\log k$, it follows that the weight of $f$ is $\Omega(k\log k)$, as desired. ∎
@@
If $S^{\prime}$ is good, then we know that, for every $u\in S_{0}^{\prime}$, there are at least $\left\lvert S_{0}^{\prime}\right\rvert/2$ vertices $v\in S_{0}^{\prime}$ with $c_{t}(u)\neq c_{t}(v)$. For any such $u,v\in S_{0}^{\prime}$ and any $w\in S_{1}^{\prime}$, it is clear that uvwuvw is rainbow. Therefore, we have found at least
[…]
@@
If we now apply Lemma 3.4 with $X$ the set of bad vertices in $T$, we see that if the total weight on the vertices in $X$ is less than $k\log k/C$, then the number of triples of leaves whose least common ancestor is outside $X$, and therefore good, is at least $k^{2.5}$. By the observation above, this would mean that we have at least $k^{2.5}/4$ rainbow triangles in $S$. We may therefore assume that the total weight on the set of bad vertices is at least $k\log k/C$. Since the weight of the vertex corresponding to $S^{\prime}$ is $\left\lvert S^{\prime}\right\rvert-2$ and the probability that $S^{\prime}$ is bad can be bounded by $\exp(-\Omega((\left\lvert S^{\prime}\right\rvert-2))\log A)$, the probability that all of the vertices in $X$ are bad is at most $\exp(-\Omega(k\log k\log A))$. Since there are at most $2^{k}$ possible choices for $X$ (as there are exactly k1k-1 internal vertices), a union bound implies that the probability there are fewer than $k^{2.5}/4$ rainbow triangles in $S$ is at most $\exp(-\Omega(k\log k\log A))$.
@@ 4 Concluding remarks
[…]
@@
Our methods prove this conjecture for $s=2$ and 4.2 follows immediately from 4.3 as $r(H,K_{n}^{(k)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any $H\in\mathcal{H}_{s}^{(k)}$. We note that all the proofs in this paper followed this strategy. Moreover, we observe that 4.3, if true in general, would also determine the tower height of diagonal hypergraph Ramsey numbers, perhaps the central open problem in hypergraph Ramsey theory. Indeed, $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any positive integers $s,k,n$ with $k>s$, so we would have that the tower height of $r(K_{n}^{(s)},K_{n}^{(s)})$ is at least s1s-1, agreeing with the longstanding upper bound.
To see the inequality $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$, suppose for the sake of contradiction that there is a red/blue-coloring of the edges of $K_{N}^{(k)}$ with $N=r(K_{n}^{(s)},K_{n}^{(s)})$ that avoids any red $H$ in $\mathcal{H}_{s}^{(k)}$. Label the $k$ parts of each of the red $s$-tight components from $1$ to $k$. Consider the auxiliary coloring of the edges of $K_{N}^{(s)}$ where an edge $e$ is red if it is contained in a red edge $\tilde{e}$ in the original coloring and $e$ contains precisely the vertices in $\tilde{e}$ from parts $1,2,\ldots,s$ in the red $s$-tight component of $\tilde{e}$; $e$ is blue otherwise. Since $\left\lvert e\right\rvert=s$, any red $\tilde{e}$ containing $e$ belongs to the same red $s$-tight component, showing that the color of $e$ does not depend on the choice of $\tilde{e}$. It is clear that each red $\tilde{e}$ in $K_{N}^{(k)}$ contains exactly one red $e$ in the auxiliary coloring of $K_{N}^{(s)}$. By the definition of $r(K_{n}^{(s)},K_{n}^{(s)})$, this auxiliary coloring contains either a red $K_{n}^{(s)}$ or a blue $K_{n}^{(s)}$. We will show that either one would give a blue $K_{n}^{(k)}$ in the original coloring. Indeed, suppose that $S$ is the set of $n$ vertices that forms a monochromatic $K_{n}^{(s)}$ in the auxiliary coloring. Then each $\tilde{e}\in\binom{S}{k}$ contains either 0 or $\binom{k}{s}>1$ red edges $e$, so it has to be blue. This yields the required contradiction and so $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ must hold.